Measures of variability :
B : 75, 85, 92, 90, 73, 83
x̅ =
(x − x̅) =
|x − x̅| =
(x − x̅)²=
Range:
R = H − L =
Mean Absolute Deviation:
MAD =
Standard Deviation:
SD =
Variance:
SD² =

Answer by Edwin McCravy(20065)

Measures of variability :
B : 75, 85, 92, 90, 73, 83

x̅ = ADD THOSE NUMBERS UP AND THEN DIVIDE BY THE NUMBER OF NUMBERS

(x − x̅) = Substitute 75 for x and what you got for x̅
|x − x̅| = Take the absolute value of the above
(x − x̅)²= Square what you just got.

(x − x̅) = Substitute 85 for x and what you got for x̅
|x − x̅| = Take the absolute value of the above
(x − x̅)²= Square what you just got.

(x − x̅) = Substitute 92 for x and what you got for x̅
|x − x̅| = Take the absolute value of the above
(x − x̅)²= Square what you just got.

(x − x̅) = Substitute 90 for x and what you got for x̅
|x − x̅| = Take the absolute value of the above
(x − x̅)²= Square what you just got.

(x − x̅) = Substitute 73 for x and what you got for x̅
|x − x̅| = Take the absolute value of the above
(x − x̅)²= Square what you just got.

(x − x̅) = Substitute 83 for x and what you got for x̅
|x − x̅| = Take the absolute value of the above
(x − x̅)²= Square what you just got.

Range:
R = H − L = The HIGHest number in B (at the top) MINUS the LEAST number in B.

Mean Absolute Deviation:
MAD = Add all the absolute values you found above and then divide by the number
of numbers.


Standard Deviation: (This was asked in the wrong place, for it's the square root
of what is asked below. So I'll answer it below the next instruction.  
SD =

Variance:
SD² = ADD all the (x − x̅)2 and then divide by either:

(a) the number of numbers if B is a population.
(b) one less than the number of numbers if B is a sample from a population.

Standard Deviation: 
SD = Take the square root of what you got for the Variance.

Edwin